### v2

# Quokka: Accelerating Program Verification with LLMs via Invariant Synthesis

Abstract
Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge. We investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants. We introduce Quokka, a first-order and effective framework for LLM-based invariant synthesis that provides sound evaluation while achieving state-of-the-art speedup results. Unlike prior work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure. We construct a benchmark of 866 instances and evaluate 9 state-of-the-art LLMs across multiple model families. Our results show that Quokka consistently outperforms all prior LLM-based verifiers: achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach. We further demonstrate that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
Machine Learning, ICML
1 Introduction
Program verification aims to provide formal guarantees that software behaves as intended, with applications in many safety-critical domains (Fan et al., 2017; Luckcuck et al., 2019). A long-standing challenge in this area, studied for more than four decades, is the automatic discovery of loop invariants. In this work, we investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants.
Loop invariants are conditions that hold before and after each loop iteration, and they are central to deductive program verification. To accelerate program verification, loop invariants must not only be correct but also sufficiently strong to prove the assertions. Generating correct invariants is relatively easy, since any universally true condition qualifies. However, only strong invariants can reduce verification effort and lead to a speedup. For example, in Figure˜1, the invariant $x>0$ is correct but not strong enough to prove the final assertion $x\neq 700$, whereas $x\equiv 3\pmod{7}$ is both correct and sufficiently strong.
Discovering such invariants is difficult and undecidable in general, which has motivated a long line of research. Traditional approaches include constraint solving (Colon et al., 2003; Gupta et al., 2009), dynamic analysis (Le et al., 2019), etc. Since invariant discovery is challenging, researchers have tried a variety of learning-based methods (Li et al., 2017; Ezudheen et al., 2018). Building on this progression, the strong capabilities of LLMs in code generation and program reasoning (Austin et al., 2021; Chen et al., 2021; Wei et al., 2025b) naturally motivate a systematic evaluation of their potential for invariant discovery.
The first work (Pei et al., 2023) to evaluate LLMs for invariant generation has two key limitations. One limitation is that it only checks correctness without assessing whether invariants are strong enough to accelerate verification. Another limitation is that correctness is determined by directly comparing against Daikon (Ernst et al., 2007), a dynamic analysis tool that infers invariants from test executions rather than formal verification. This approach is unsound: Daikon’s invariants may themselves be incorrect. Moreover, semantically equivalent invariants (e.g., $a>0$ vs. $a\geq 1$ for integers) can be incorrectly rejected as wrong. Consequently, prior work cannot reliably evaluate either the correctness or the practical utility of LLM-generated invariants.
A series of follow-up works have proposed LLM-based verifiers. Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs. This line of work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants. For instance, LaM4Inv (Wu et al., 2024a) uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through complex logical operations (conjunctions or disjunctions) to construct valid loop invariants, rather than using them directly. LOOPY (Kamath et al., 2024) employs a customized Houdini algorithm with iterative refinement to filter candidate invariants. LEMUR (Wu et al., 2024b) develops a backtracking and sequential algorithm. While these approaches demonstrate the potential of LLM-based verification, their complexity raises a question: can we design a simpler, principled, and more effective decision procedure for LLM-based verifiers?
Figure 1: Illustration of Quokka’s evaluation pipeline. The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3). The verification procedure then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
In this work, we introduce Quokka, a framework for LLM-based invariant synthesis that accelerates program verification. Quokka provides a sound approach to evaluate LLMs’ capabilities in generating invariants while achieving state-of-the-art speedup results compared with prior LLM-based verifiers. We formalize our methodology as a proof calculus, providing a rigorous foundation for building LLM-based verifiers. Instead of checking the correctness of invariants based on string-matching, we adopt a verifier-based approach that directly queries the underlying solver. Unlike previous work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure. Our approach enables both soundness guarantees and precise measurement of performance speedup, allowing us to systematically evaluate whether LLM-generated invariants can accelerate verification in practice.
To support systematic comparison across solvers and LLMs, we construct a dataset of 866 instances derived from the most recent edition of the software verification competition SV-COMP (Beyer and Strejček, 2025). To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date, compared with datasets used in previous work. Quokka consistently outperforms all prior LLM-based verifiers. When comparing against LEMUR (the previous best LLM-based verifier with the same underlying solver), Quokka achieves speedups of at least 1.2× (i.e., verification completes at least 20% faster than the baseline solver) on 81 instances compared to LEMUR’s 39 instances. For more substantial speedups of at least 2.0× (i.e., verification completes at least twice as fast), Quokka reaches 51 instances compared to LEMUR’s 22 instances.
We evaluate 9 state-of-the-art LLMs, including GPT models from OpenAI, Claude models from Anthropic, and models from the Qwen and LLaMA families. Our experiments reveal that gpt-5.2 achieves the strongest speedup results among all evaluated models. Beyond model evaluation, we investigate techniques to enhance invariant generation capabilities. We construct a training dataset of 3589 instances using a verifier-based filtering approach that ensures data quality. Our experiments demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification. Specifically, Best-of-8 sampling achieves speedups of at least 1.2× on 99 instances, representing a 22% improvement over single-sample generation, which achieves such speedups on only 81 instances.
In summary, our contributions are as follows:
  • We introduce a decision procedure for LLM-based invariant synthesis that provides sound evaluation and achieves the best speedup results compared with prior LLM-based verifiers.
  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 state-of-the-art LLMs across multiple model families.
  • We construct training data and demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
2 Related Work
Traditional Methods for Program Invariant Generation.
A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan and Qadeer, 2002; Lahiri and Bryant, 2007; Hojjat and Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot and Cousot, 1977; Cousot and Halbwachs, 1978; Cousot and Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala and McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich and Bodík, 2018). Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019). Daikon executes the program, observes runtime values, and reports properties that consistently hold over the observed executions. However, such invariants may fail to generalize to all possible executions, thereby compromising soundness. Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
Learning-Based Method for Invariant Generation.
Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014, 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020). More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a, b, c), giving rise to a series of work that explore using LLMs for finding invariants. (Pei et al., 2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation. Various techniques have been proposed to couple LLMs with symbolic solvers, including ranking LLM-generated invariants (Chakraborty et al., 2023), the “query-filter-reassemble” strategy of LaM4Inv (Wu et al., 2024a), the back-tracking algorithm in LEMUR (Wu et al., 2024b), and Loopy’s integration of the classic Houdini algorithm (Kamath et al., 2024). On the dataset side, (Liu et al., 2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach. In contrast, our work provides a simple and sound evaluation procedure for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
3 Method
3.1 Preliminary
We formalize the task of loop invariant synthesis using standard Hoare logic (Hoare, 1969). A Hoare triple $\{P\}\,S\,\{Q\}$ specifies that if the precondition $P$ holds before executing a statement $S$, then the postcondition $Q$ will hold after its execution. In the context of loops, an invariant $I$ is a logical proposition that summarizes the state of the program at each iteration, and it is the key to proving the validity of Hoare triples involving loops. For a loop of the form while $B$ do $S$, the goal of invariant synthesis is to identify a loop invariant $I$ that satisfies the following inference rule:
$\frac{P\Rightarrow I\quad\{I\wedge B\}\,S\,\{I\}\quad I\wedge\lnot B\Rightarrow Q}{\{P\}\ \texttt{while}\ B\ \texttt{do}\ S\ \{Q\}}$
Here, $P$ is the precondition, $Q$ is the postcondition, $B$ is the loop condition, and $S$ is the loop body. Intuitively, the inference rule requires that the invariant $I$ holds at the beginning of the loop ($P\Rightarrow I$), is preserved across every iteration of the loop body ($\{I\wedge B\}\,S\,\{I\}$), and upon termination ensures the postcondition ($I\wedge\lnot B\Rightarrow Q$).
Invariant synthesis amounts to generating a logical summary $I$ that is both correct, meaning it can be verified, and strong, meaning it enables verification of the final assertion. Weak but correct invariants contribute little, leaving most of the reasoning to the verifier, whereas strong invariants narrow the search space of program states, reduce solver effort, and yield substantial speedups.
3.2 Verifier-Based Decision Procedure
We formalize our verifier-based procedure for assessing candidate invariants. Let $P$ denote a program. A property is written as $p=\langle\varphi,\ell\rangle$, where $\varphi$ is a state predicate and $\ell$ is a program location. For a finite set $A$ of properties, let $\mathrm{Asm}(P,A)$ be the program obtained from $P$ by inserting assume statements for all elements of $A$. An execution of $\mathrm{Asm}(P,A)$ that reaches a location where an assumption is violated terminates immediately. We write $P\models_{A}p$ to indicate that all executions of $\mathrm{Asm}(P,A)$ satisfy the assertion $p$. The notation $P\models p$ abbreviates $P\models_{\varnothing}p$. Since assumptions restrict behaviors, if $P\not\models_{A}p$ for some $A$, then necessarily $P\not\models p$.
We assume access to a verifier
$V(P,A,p)\in\{\mathbf{T},\mathbf{F},\mathbf{U}\},$
which returns either $\mathbf{T}$ (proved), $\mathbf{F}$ (refuted), or $\mathbf{U}$ (inconclusive). The verifier is required to be sound on conclusive outcomes:
$V(P,A,p)=\mathbf{T}\;\Rightarrow\;P\models_{A}p,$
$V(P,A,p)=\mathbf{F}\;\Rightarrow\;P\not\models_{A}p.$
No completeness is assumed for $\mathbf{U}$, which may arise from timeouts or incompleteness of the underlying verifier.
The verification task specifies a target property $p^{\star}=\langle\varphi^{\star},\ell^{\star}\rangle$. Given $P$ and $p^{\star}$, a large language model proposes a candidate invariant $q=\langle\psi,\ell\rangle$, typically at a loop header. To evaluate the utility of $q$, the procedure issues two verifier queries:
$d_{a}:=V(P,\varnothing,q)$
$d_{b}:=V(P,\{q\},p^{\star})$
The query $d_{a}$ verifies whether $q$ is a correct invariant, while the query $d_{b}$ determines whether the target property holds under the assumption that $q$ is true. An illustration of these two verifier queries is provided in Figure˜1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure). The outcome of the procedure is expressed as a judgment
$P\;\Rightarrow\;\langle p^{\star},q\rangle\;\Downarrow\;d\qquad\text{with}\quad d\in\{\mathbf{T},\mathbf{F},\mathbf{U}\}.$
The interpretation is as follows: if the judgment yields $\mathbf{T}$, then $p^{\star}$ is established on $P$; if it yields $\mathbf{F}$, then $p^{\star}$ is refuted; and if it yields $\mathbf{U}$, the attempt is inconclusive.
The inference rules defining this judgment are given below. Each rule specifies one possible derivation of the outcome, depending only on the responses of the verifier.
[…]
Rule DEC-FALSE captures short-circuit refutation: if the goal fails even in the restricted program $\mathrm{Asm}(P,\{q\})$, then it is false on the original program $P$. Rule DEC-PROP implements the prove-then-use strategy: once the candidate invariant $q$ is established, the outcome is exactly the verifier’s answer on the goal under the assumption $q$, restricted to $d\in\{\mathbf{T},\mathbf{U}\}$ so as not to overlap with DEC-FALSE. Rule DEC-$\mathbf{U}$ gives explicit conditions for inconclusiveness: the goal is not refuted under $q$ and $q$ is not established as an invariant.
Theorem (Decision Soundness).
If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{T}$ is derivable, then $P\models p^{\star}$. If $P\Rightarrow\langle p^{\star},q\rangle\Downarrow\mathbf{F}$ is derivable, then $P\not\models p^{\star}$.
Proof.
For the $\mathbf{T}$ case, the final rule must be DEC-PROP with $d=\mathbf{T}$. The premises yield $V(P,\varnothing,q)=\mathbf{T}$ and $V(P,\{q\},p^{\star})=\mathbf{T}$. By verifier soundness, $P\models q$ and $P\models_{\{q\}}p^{\star}$. Since $q$ holds on all executions of $P$, introducing the assumption $\{q\}$ does not remove executions relevant to $p^{\star}$; thus $P\models p^{\star}$. For the $\mathbf{F}$ case, the final rule must be DEC-FALSE. Its premise $V(P,\{q\},p^{\star})=\mathbf{F}$ implies $P\not\models_{\{q\}}p^{\star}$ by soundness. Assumptions restrict behaviors; hence, a violation under assumptions entails a violation without them, yielding $P\not\models p^{\star}$. ∎
This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct. The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
3.3 Implementation
We describe the implementation of our verifier-based evaluation framework. Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the decision procedure. Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
Syntactic Validation.
Before invoking the verifier, we apply syntactic checks to the generated predicate $\psi$. These checks ensure that $\psi$ can be safely interpreted as a state predicate and that its insertion as an assumption does not alter the program state. For instance, expressions that update variables (e.g., a += 1) are rejected. Only Boolean conditions over the program state are accepted.
Parallel Verifier Queries. For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$. These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification. The final outcome is then derived exactly according to the decision calculus.
3.4 Supervised Fine-Tuning and Best-of-N Sampling
Below, we discuss how we construct our dataset for fine-tuning and the way we perform Best-of-N sampling.
Synthetic Dataset Generation.
To construct the synthetic dataset, we prompt GPT-4o using the template in Appendix A.1. The template takes three seed programs as examples and instructs the model to synthesize a new C program that is compilable and contains both loops and assertions. To obtain a diverse and large pool of candidates, we repeatedly invoke the model with different seed programs. To avoid data leakage, these seed programs are randomly drawn from the SV-COMP pool  (Beyer and Strejček, 2025) that is disjoint from our evaluation set. Although the prompt requests compilable programs with loops and assertions, the LLM-generated programs may fail to compile, include assertions that do not hold, and contain multiple assertions. For programs with multiple assertions, we split them into separate instances, each retaining only a single assertion while preserving all loop structures (ensuring at least one loop per instance). We then run UAutomizer on every program and discard any instance that is non-compilable or whose assertion is invalid. This filtering step ensures the quality of the dataset, resulting in 3589 synthetic programs.
Extract Invariants Generated from UAutomizer.
When running UAutomizer to prove the assertions in the synthetic programs, the tool also emits the invariants it discovers. From its output, we extract the loop invariants. Each extracted invariant includes its program location and its predicate. Although each program contains exactly one assertion, it may include multiple loops, so a single program can yield multiple loop invariants, all associated with the same assertion. We pair each program with each of its corresponding loop invariants to form our training dataset. We show an example invariant generated from UAutomizer in Appendix A.2. With the ground-truth invariants, we perform supervised fine-tuning using LoRA (Hu et al., 2022).
Best-of-N Sampling.
Best-of-N sampling is an inference-time strategy in which multiple candidate programs are generated, and the most effective one is selected, a technique shown to improve performance on code-generation tasks (Ehrlich et al., 2025). In our setting, the best candidate is the invariant that yields the largest speedup, i.e., the one whose decision procedure finishes earliest. As described in Section˜3.2, evaluating a single candidate requires two verifier queries issued in parallel; therefore, Best-of-N sampling evaluates 2N verifier queries concurrently.
4 Experimental Setup
Dataset.
We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis. We collect a pool of 866 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record per-instance solving times. The process is fully automated, with no manual filtering or cherry-picking. We use this timeout to enable meaningful measurement of verification speedups. In contrast to prior work (Wu et al., 2024b), which evaluates only instances that the baseline cannot solve within this time budget, our approach requires knowing the baseline runtime to quantify relative improvements. Restricting to instances whose baseline runtime is observable allows us to report concrete speedup factors, rather than only counts of additionally solved problems. To the best of our knowledge, this dataset is the largest used in prior evaluations of LLM-based verifiers. Also, the programs in our dataset are significantly longer, with an average of 53 lines of code, compared to only 22 for LaM4Inv (Wu et al., 2024a), 23 for LEMUR (Wu et al., 2024b), and 27 for Loopy (Kamath et al., 2023). Manual inspection shows that our dataset contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets. This makes Quokka a more challenging and realistic benchmark that better distinguishes solver performance.
For training data, we construct our training data with synthetic data generation filtered by the solvers, as described in Section˜3.4. We make sure that the seed programs used for synthetic data generation are disjoint from our evaluation dataset, mitigating the leakage issue. The statistics of our dataset are shown in Table˜1.
Dataset Avg. #Lines #Instances
Evaluation 53 866
Training 42 3589
Table 1: Dataset statistics.
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
Qwen2.5-7B 419 34 26 23 20 18
o3 550 38 33 30 24 23
Llama-3.1-8B 342 43 35 32 29 27
claude-opus-4.1 487 56 45 38 35 31
Qwen2.5-72B 500 63 49 45 37 33
claude-opus-4.5 689 69 53 46 39 34
gpt-5.1 694 56 46 41 40 37
claude-sonnet-4 620 68 58 51 45 41
gpt-5.2 710 81 65 60 56 51
Table 2: Performance of different LLMs when using Quokka. We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
Metrics.
We evaluate LLMs along two dimensions: (1) the correctness of the generated invariants and (2) the performance improvements they provide. Correctness is determined by the solver using a timeout set to $1.2\times$ the original solving time for each problem. This slack allows the solver additional time to assess the correctness of model-generated invariants, reducing inconclusive outcomes caused by timeouts. All speedup-related measurements are reported relative to the original solver, which serves as the baseline. We report speedup using the metric of $\mathtt{\#Fast}_{p}$, defined as the number of instances for which verification completes at least $p\times$ faster than the baseline, with $p\in\{1.2,1.4,\ldots\}$. Additionally, we also report the number of instances solved under varying timeout budgets.
Choice of Solvers.
We use the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024) as our base solver. Our methodology requires that the solver can verify programs without externally-provided invariants, enabling us to measure speedup relative to a functional baseline. While prior work has explored other solvers such as Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018), we select UAutomizer for two key reasons. First, UAutomizer represents the current state-of-the-art in automated verification, outperforming ESBMC on standard benchmarks in SV-COMP. Second, unlike Frama-C, which requires externally-provided invariants and cannot proceed without them, UAutomizer has built-in invariant synthesis capabilities that allow it to verify programs autonomously. To demonstrate the generality of Quokka’s methodology across different solvers, we also present experiments with ESBMC in Appendix A.5, where we show that Quokka successfully accelerates ESBMC as well.
Models.
We benchmark Claude models from Anthropic, GPT models from OpenAI, and models from the Qwen and LLaMA families (Hui et al., 2024; Yang et al., 2025; Touvron et al., 2023).
Hardware and OS.
Experiments were conducted on a node with Intel E5-2640 v4 CPUs, 256G main memory, running Ubuntu 20.04 LTS.
5 Results
5.1 Results of LLMs
Table˜2 presents the performance of different LLMs when using Quokka. The results are sorted by the number of instances achieving at least $2.0\times$ speedup (# Fast2.0), which ranges from 18 instances for Qwen2.5-7B to 51 instances for gpt-5.2. Overall, the table demonstrates substantial variation in both correctness and speedup performance across different models, with larger and more recent models generally achieving better results.
Effectiveness of Quokka.
Our approach demonstrates the speedup capabilities across all evaluated models. Even smaller models like Qwen2.5-7B achieve speedups on 34 instances at the $1.2\times$ threshold, while the best-performing model gpt-5.2 accelerates 81 instances at this level and achieves $2.0\times$ or greater speedup on 51 instances. These results demonstrate that Quokka can effectively leverage LLMs to generate invariants that meaningfully accelerate verification. The consistent speedup gains across different model families, from 7B parameter models to frontier models, and the progressive improvement with model scale underscore the effectiveness of our approach.
Correctness vs. Performance Gap.
A striking observation is the substantial gap between generating correct invariants and achieving meaningful speedups. For example, gpt-5.2 produces 710 correct invariants but achieves $2.0\times$ speedup on only 51 instances, while claude-opus-4.5 generates 689 correct invariants yet delivers $2.0\times$ speedup on merely 34 instances. This gap highlights that correctness alone is insufficient: invariants must also be strong enough to meaningfully accelerate verification, requiring not only logical correctness but also strategic insight into which invariants will most effectively guide the solver.
Model Inference Time.
We include model inference time in our evaluation to provide a fair and realistic assessment of end-to-end performance. As documented in Appendix A.4, smaller models such as Llama-3.1-8B and Qwen2.5-7B exhibit negligible generation times (0.3–0.4 seconds), while most models complete inference within 3 seconds. The o3 model has the highest overhead at 8.1 seconds on average. Importantly, unlike LOOPY (Kamath et al., 2023), which invokes the LLM up to 15 times per program and thus spends a substantial portion of its time budget on API calls, our approach requires only a single LLM invocation per program. Consequently, LLM inference accounts for only a small fraction of total runtime. Furthermore, our evaluation dataset contains sufficiently challenging programs (with solving times up to 600 seconds) such that solver time naturally dominates inference overhead, making the comparison meaningful even when inference latency is included.
Prompting-Based Methods.
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As detailed in Appendix A.6, we evaluate 0-shot, 2-shot, 4-shot, and chain-of-thought (CoT) prompting. Few-shot prompting does not consistently improve performance, while CoT prompting yields the weakest results, likely due to increased generation time without proportional quality gains. These findings indicate that simple zero-shot prompting is most effective for invariant synthesis.
5.2 Comparison with Prior LLM-based Verifiers
Method # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
LaM4Inv 1 1 1 1 1
LOOPY 8 7 5 3 1
LEMUR-original 19 16 13 12 10
LEMUR-gpt-5.2 39 30 28 25 22
Quokka-gpt-5.2 81 65 60 56 51
Table 3: Comparison of speedup performance across different methods. We report the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
Table˜3 compares the speedup performance of Quokka against prior LLM-based verifiers across different speedup thresholds, while Figure˜2 shows the number of solved instances over time for each method.
Quokka Outperforms Prior LLM-based Verifiers.
Quokka with gpt-5.2 achieves state-of-the-art results across all speedup metrics. Among prior work, LEMUR (Wu et al., 2024b) represents the strongest baseline. To ensure a fair comparison, we enhance LEMUR by replacing its original model (gpt-4) with gpt-5.2, creating LEMUR-gpt-5.2. Even with this enhancement, Quokka-gpt-5.2 significantly outperforms LEMUR-gpt-5.2: at the $1.2\times$ speedup threshold, Quokka-gpt-5.2 accelerates 81 instances versus 39 for LEMUR-gpt-5.2 and 19 for LEMUR-original. At $2.0\times$ speedup, Quokka-gpt-5.2 achieves 51 instances compared to 22 for LEMUR-gpt-5.2 and 10 for LEMUR-original. LaM4Inv (Wu et al., 2024a) and LOOPY (Kamath et al., 2023) demonstrate minimal speedup capabilities, achieving at most 1 and 8 instances respectively at the $1.2\times$ level. These results demonstrate that Quokka establishes a new state-of-the-art for LLM-based verification acceleration.
Differences in Base Solvers.
Different LLM-based verification frameworks are built on different base solvers: Quokka and LEMUR use UAutomizer, LOOPY uses Frama-C (Cuoq et al., 2012), and LaM4Inv uses ESBMC (Gadelha et al., 2018). The choice of base solver significantly impacts the performance, as stronger solvers provide a better foundation for speedup. Since UAutomizer represents the current state-of-the-art solver, we believe future work should prioritize developing LLM-based verifiers atop state-of-the-art solvers such as UAutomizer. For a direct comparison between UAutomizer and prior LLM-based solvers, we also refer readers to Appendix A.3, where we evaluate UAutomizer on the dataset released by each of the prior works.
Algorithmic Simplicity and Parallelism.
Prior work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants. In contrast, Quokka employs a simple, verifier-based, and parallel approach that directly leverages the base solver’s capabilities.
LaM4Inv uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through a complex algorithm, rather than using them directly. LOOPY employs a highly-customized Houdini algorithm with iterative fixing to filter candidates. LEMUR is closest to our approach, also directly using the base solver to verify the correctness of invariants, but follows a purely sequential decision process with backtracking. In contrast, Quokka directly uses LLM-generated invariants without decomposition or reassembly, and exploits parallelism by issuing the verification queries in parallel. This design reduces implementation complexity while enabling more efficient hardware utilization, contributing to the superior speedup performance demonstrated in our experiments.
Figure 2: Number of instances solved by different methods over varying timeout budgets.
Model # Incorrect Invariants # Goal Timeout # Invariant Timeout # Both Timeout
gpt-5.1 78 96 245 110
claude-sonnet-4 110 160 209 145
gpt-5.2 56 139 230 121
Table 4: Failure mode breakdown for the top three models on Quokka. Incorrect Invariants: the invariant is refuted by the solver. Goal Timeout: the invariant is verified, but proving the assertion under it times out. Invariant Timeout: verifying the invariant itself times out. Both Timeout: both verification queries time out.
Model # Correct Invariants # Fast1.0 # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
Qwen2.5-7B (base) 419 116 34 26 23 20 18
Qwen2.5-7B (finetuned) 431 141 39 27 26 24 20
Table 5: Performance comparison of base and fine-tuned Qwen2.5-7B. Fine-tuning leads to modest improvements in both correctness and speedup metrics.
5.3 Failure Mode Analysis
To understand why many invariants fail to accelerate verification in Quokka, we analyze failure modes for the top three models. We categorize failures into four types: 1) Incorrect Invariants: the invariant is refuted by the solver; 2) Goal Timeout: the invariant is verified, but proving the assertion under it times out; 3) Invariant Timeout: verifying the invariant itself times out; 4) Both Timeout: both verification queries time out. The results in Table˜4 reveal that while some failures stem from incorrect invariants, the dominant issue is timeout-related. This suggests that current models lack an understanding of what makes verification queries tractable for symbolic solvers. Future work should explore training reward models or other mechanisms to help models internalize solver difficulty, enabling them to propose invariants that genuinely simplify verification by decomposing hard problems into easier subproblems rather than inadvertently generating verification conditions that are as difficult or harder to solve than the original proof goal.
Figure 3: Effect of Best-of-N sampling on the number of instances under different Fastp settings.
5.4 Results of Fine-Tuning and Best-of-N Sampling
Modest Improvement from Supervised Fine-Tuning.
We perform supervised fine-tuning using LoRA (Hu et al., 2022) on Qwen2.5-7B for 3 epochs. As shown in Table˜5, fine-tuning yields modest improvements: correct invariants increase from 419 to 431, and instances achieving meaningful speedups improve across all thresholds. For example, Fast1.0 improves from 116 to 141, and Fast1.2 improves from 34 to 39. These results suggest that domain-specific fine-tuning can improve both correctness and speedup, though gains remain moderate.
Benfits of Best-of-N Sampling.
We evaluate Best-of-N sampling by generating multiple invariant candidates from gpt-5.2 at temperature 0.7 and selecting the one yielding the fastest verification time. As shown in Figure˜3, increasing N from 1 to 4 consistently improves speedup across all Fastp thresholds. However, performance plateaus beyond N=8 due to hardware constraints: our 20-core CPU must verify 2N queries in parallel, and when N exceeds 10, resource contention may negate sampling benefits. Memory limits can further constrain scalability. While Best-of-N sampling is effective, its benefits are hardware-bounded, though modest values of N yield meaningful improvements.
6 Conclusion
This work introduced Quokka, a simple yet effective framework for LLM-based invariant synthesis that accelerates program verification. Our approach employs a principled verification procedure that decomposes the original proof goal into two independent subproblems and launches verifier queries in parallel, providing sound evaluation while achieving state-of-the-art speedup results. Using a benchmark of 866 instances, we evaluated 9 state-of-the-art LLMs across multiple model families and compared against prior LLM-based verifiers. The results demonstrate that Quokka consistently outperforms all existing LLM-based verifiers, achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach. We further demonstrated that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification. Our work establishes a new state-of-the-art for LLM-based program verification and provides a principled foundation for future research in this direction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Prompt Template Used for Training Dataset Generation
Table˜A1 shows the prompt template used for synthesizing training programs from seed programs.
You will be shown 3 example C programs. Please gain inspiration from the following programs to create a new high-quality C program. Do not simply copy from any of them. Requirements for the generated program: 1. The program MUST contain non-trivial loops (for or while). 2. The program MUST contain assertions. 3. The program MUST be compilable, self-contained, and reasonably complex (not trivial or overly short). 4. Only output the new C program. Example snippets: Program 1: {SEED_PROGRAM_1}
Program 2:
{SEED_PROGRAM_2}
Program 3:
{SEED_PROGRAM_3}
Output format: The generated program must be wrapped strictly in the following format: ‘‘‘c <NEW_C_PROGRAM>
Table A1: Prompt template for synthetic data generation.
A.2 An Example from the Fine-Tuning Dataset
Original Program (Input to UAutomizer)
1int main() {
2 int N = __VERIFIER_nondet_int();
3 assume_abort_if_not(N >= 1 && N < 100);
4 int x = 0, y = 0;
5 for (int i = 0; i < N; i++) {
6 if (i % 2 == 0) {
7 x += i;
8 } else {
9 y += i;
10 }
11 }
12 int diff = x - y;
13 __VERIFIER_assert(diff <= N);
14 return 0;
15}
Loop Invariant Generated by UAutomizer:
1Line Number: 6
2Predicate:
3(
4(i + 1) % 2 == 0 &&
5x < 2 + i + y &&
6x < 2 + y + 2 * i &&
7x < N + y + 1 &&
81 <= N
9)
10||
11(
12i % 2 == 0 &&
13x < 2 + i + y &&
14x < 2 + y &&
151 <= N
16)
Figure A1: An example from the fine-tuning dataset: program and its loop invariant generated by UAutomizer.
Figure˜A1 shows an example from the fine-tuning dataset with the program to the UAutomizer and the generated loop invariant.
The loop invariant holds at Line 6 (the beginning of the loop). It is a disjunction of two clauses, and can be written as $I\equiv P\lor Q=\big((i+1)\bmod 2=0\land x<2+i+y\land x<2+y+2i\land x<N+y+1\land 1\leq N\big)\lor\big(i\bmod 2=0\land x<2+i+y\land x<2+y\land 1\leq N\big)$.
By inspecting the loop, we can derive an exact relationship between $x$, $y$, and $i$ at the beginning of each iteration. Since $x$ accumulates all even numbers less than $i$ and $y$ accumulates all odd numbers less than $i$, we obtain:
$\text{if $i$ is even:}\quad x-y=-\frac{i}{2},\qquad\text{if $i$ is odd:}\quad x-y=\frac{i-1}{2}.$
Using this relationship, it is straightforward to verify that the invariant $I$ produced by UAutomizer is correct. Moreover, $I$ is strong enough to prove the final assertion $x-y\leq N$.
To show that the assertion holds at loop termination, we can check the following verification condition
$I\;\wedge\;(i=N)\;\Rightarrow\;(x-y\leq N),$
where $i=N$ denotes the loop’s exit condition.
Case 1: $N$ is odd.
Instantiating the invariant with $i=N$ activates the $P$ disjunct of $I$, from which we obtain
xy<N+1.x-y<N+1.
Since xyx-y is an integer, this directly implies $x-y\leq N$.
Case 2: $N$ is even.
In this case, the $Q$ disjunct applies. From the clause $x<2+y$ contained in $Q$, we derive
xy<2.x-y<2.
Because xyx-y is an integer and $N\geq 1$ and even (hence $N\geq 2$), we conclude
$x-y\leq 1\leq N,$
establishing the desired post-condition.
Thus, the invariant $I$ indeed suffices to prove the final assertion.
A.3 Results of UAutomizer on Prior Datasets
Verifier Total Instances Solved Instances under Different Timeouts
10s 100s 300s 600s
LaM4Inv (Wu et al., 2024a) 316 144 286 295 299
UAutomizer 316 299 299 299 299
Loopy (Kamath et al., 2023) 469 0 133 353 403
UAutomizer 469 372 403 411 413
LEMUR (Wu et al., 2024b) 47 2 8 16 19
UAutomizer 47 0 0 0 0
Table A2: Comparison of prior LLM-based verifiers and UAutomizer on their own custom dataset under different timeout budgets.
As shown in Table A2, UAutomizer, the state-of-the-art non-LLM-based verifier, consistently solves more instances than LaM4Inv (Wu et al., 2024a) and Loopy (Kamath et al., 2023), two representative LLM-based tools, on their respective custom datasets. This demonstrates the importance of base solver choice.
LEMUR (Wu et al., 2024b) is the only tool that surpasses UAutomizer on its own benchmark. This outcome is largely explained by LEMUR’s dataset construction methodology: it specifically curates problems that UAutomizer cannot solve within 600 seconds. To assess whether LEMUR’s advantage generalizes beyond this carefully selected benchmark, we conduct an additional analysis. We randomly selected 50 instances that UAutomizer failed to solve within 600 seconds and evaluated LEMUR’s performance on them. LEMUR successfully solved 12 of these 50 instances within the same 600-second timeout, demonstrating that its gains are not merely an artifact of benchmark design but reflect genuine capability on challenging problems. Nevertheless, for our work, we focus on problems that UAutomizer can solve within the time budget, as our primary objective is to quantify speedup.
A.4 Inference Time of Different Models
Table˜A3 presents the average inference time per instance for different LLMs when generating invariants. As expected, smaller open-source models such as Llama-3.1-8B and Qwen2.5-7B exhibit the fastest generation times, averaging around 0.3–0.4 seconds per instance. Mid-sized models like Qwen2.5-72B and the GPT-5 series require approximately 0.6–1.4 seconds. The Claude models take longer at around 1.6–2.7 seconds per instance. The o3 model has the highest inference overhead, averaging over 8 seconds per instance. These results highlight the trade-off between model capability and inference cost, which must be considered when evaluating end-to-end verification performance.
Model Avg Generation Time (s)
Llama-3.1-8B 0.300
Qwen2.5-7B 0.355
Qwen2.5-72B 0.578
gpt-5.2 1.076
gpt-5.1 1.373
claude-opus-4.1 1.626
claude-opus-4.5 2.538
claude-sonnet-4 2.746
o3 8.075
Table A3: Average inference time per instance for different LLMs when generating invariants.
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
gpt-5.1 399 10 9 9 9 9
gpt-5.2 455 17 11 10 9 9
Table A4: Performance of different LLMs when using ESBMC as the base verifier. We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over ESBMC.
A.5 Results for ESBMC as the Base Solver
To demonstrate the generality of Quokka’s methodology, we conduct additional experiments using ESBMC (Gadelha et al., 2018) as the base solver instead of UAutomizer. As shown in Table˜A4, Quokka successfully accelerates ESBMC as well, with both gpt-5.1 and gpt-5.2 generating hundreds of correct invariants and achieving speedups on multiple instances. For example, gpt-5.2 produces 455 correct invariants and achieves speedups of at least $1.2\times$ on 17 instances. These results confirm that Quokka is not limited to a specific solver and can be applied to accelerate different verification tools.
A.6 Results for Prompting-Based Methods
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
gpt-5.2 (CoT) 639 41 32 26 22 20
gpt-5.2 (4-shot) 756 72 62 53 47 43
gpt-5.2 (2-shot) 764 83 67 58 53 49
gpt-5.2 710 81 65 60 56 51
Table A5: Comparison of different prompting strategies for gpt-5.2.
We investigate the impact of different prompting strategies on invariant generation quality using gpt-5.2. As shown in Table˜A5, we compare chain-of-thought (CoT) prompting, few-shot prompting with 2 and 4 examples, and our default zero-shot approach. The results show that few-shot prompting does not necessarily improve performance and may introduce noise or distract the model from the task at hand. Chain-of-thought prompting yields the weakest results across all metrics, with only 639 correct invariants and 41 instances achieving $1.2\times$ speedup. This suggests that explicit reasoning steps may increase generation time without proportional benefits to invariant quality, thereby reducing overall speedup. These findings indicate that prompting-based strategy can hardly improve the task of invariant synthesis.